OpenStudy (kingj):

please, help setting up this problem: the measure of an angle is 20 less than 3 times its supplement. find the measure of the angle.

OpenStudy (ganpat):

let angle be x and its supplement is 180 - x so, x = 3 (180-x) - 20 x = 540 - 3x - 20 4x = 520 x = 130 degrees...

OpenStudy (kingj):

thank you

OpenStudy (ganpat):

wc